1. Create a “formula” for the process you used above, such that you’d be able to determine the field of view for a lens of any magnification if you knew the field of view of another lens magnification on the same microscope.
  2. Using your answers from the previous two questions, estimate the size -- in microns -- of an organism if it covered one-third of the field of view at 100X total magnification. Show your work/process.
1
Expert's answer
2021-01-25T08:11:24-0500

1.To calculate the microscope’s field of view, divide the field number by the magnification number;


FOV ="\frac{Field Number} {Magnification Number}"



2."Field Of View =\frac{\frac{1} {3}}{1000}"


="\frac{1000} {3}"


=333.33 microns.